OpenStudy (anonymous):

Alternating series help: \[1- 1/2 +1/3 + 1/4 ... (-1)^{n+1}/n\] My book says this is an alternating series, but I thought the signs had to alternate. Is this a mistake?

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@Zarkon @UnkleRhaukus

OpenStudy (unklerhaukus):

\[\sum\limits_1^n\frac{(-1)^{n+1}}n=1- 1/2 +1/3 -1/4 +\cdots (-1)^{n+1}/n\]

OpenStudy (anonymous):

So, the book is incorrect, right?

OpenStudy (unklerhaukus):

if the book has a + in front of the quater then it is incorrect \[\frac{(-1)^{4+1}}{4}=-1/4\]

OpenStudy (anonymous):

Thank you